OpenStudy (anonymous):

Evaluate the expression. r = <9, -7, -8>, v = <3, 4, 7>, w = <6, -9, 7> v ⋅ w PLEASE HELP ME! Ive been on forever and nobody has been able to help

OpenStudy (loser66):

you have formula to calculate dot product, just apply. v.w = (3*6)+(4*(-9)+(7*7) = some number. I saw this question once but I ignore because I think it's not a problem , you just apply formula, if you don't know that formula, look at mine, or look up at your note. i'm pretty sure that you have many questions similar those stuffs . try to figure out by yourself

OpenStudy (anonymous):

So its 31?

OpenStudy (loser66):

@zepdrix why do you stop?

zepdrix (zepdrix):

Yah 31 sounds right! :) Oh I just thought it was strange that they gave you some vector \(\large \vec r\) and you don't even use it. Was this information given for a series of questions I guess? :o Or maybe the \(\large \vec r\) was just to confuse you :3

OpenStudy (anonymous):

It was only one question . I was given the choices of ' <-27, 28, 56> <18, -36, 49>, 18 or 31 and when i calculated it i got 31. so thats correct?

zepdrix (zepdrix):

Yes that's correct c:

OpenStudy (anonymous):

ok thanks do you think you could help me with another?

OpenStudy (anonymous):

@zepdrix ?

zepdrix (zepdrix):

Sure.

OpenStudy (anonymous):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. C = 37°, a = 19, c = 8 A = 90.6°, B = 52.4°, b ≈ 14 A = 52.4°, B = 90.6°, b ≈ 14 The triangle cannot be solved with the Law of Sines. No triangle is formed

OpenStudy (anonymous):

could you explain how i am supposed to do this because i get confused when trying to use the law of sines

OpenStudy (anonymous):

@zepdrix ?

zepdrix (zepdrix):

So is the part that's given to us `C = 37°, a = 19, c = 8`, While the rest is multiple choice answers?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Ok gimme a sec c:

zepdrix (zepdrix):

So let's draw a triangle just for fun.

zepdrix (zepdrix):



OpenStudy (anonymous):

fancy

zepdrix (zepdrix):

Let's try to use the Law of Sines and see what happens. \(\huge \dfrac{\color{orangered}{a}}{sin A}=\dfrac{\color{orangered}{c}}{sin \color{orangered}{C}}\) The Law of Sines tells us that we can relate the sides and angles in the given way. The reason we're not involving the B's yet is because we were GIVEN these orange terms. Making it possible to solve for the 4th one in our equation.

OpenStudy (anonymous):

so 19/sinA =8/sin37

zepdrix (zepdrix):

Ok good, let's solve for sinA from here.

OpenStudy (anonymous):

how?

zepdrix (zepdrix):

Multiply both sides by sin A, giving us, \[\large 19=\frac{8}{\sin 37}\sin A\] Then we'll multiply both sides by sin 37,\[\large 19\sin 37=8\sin A\]Then finally, divide both sides by 8,\[\large \sin A=\frac{19\sin37}{8}\] We had to jimmy some things around, but we were able to do it.

OpenStudy (anonymous):

ohh okay!! so A is -1.5

OpenStudy (anonymous):

Hold on -1.5 that mean that no triangle happens?

zepdrix (zepdrix):

Hold on! Simmer down now! We still have some work to do! :)

OpenStudy (anonymous):

oh ok . sorry

zepdrix (zepdrix):

If you punch all that into your calculator, you should come up with,\[\large \sin A=1.43\] But we ultmiately want to solve for \(\large A\), not \(\large \sin A\). So we'll rewrite this using the Inverse Sine function.\[\large \sin A=1.43 \qquad \rightarrow \qquad A=\arcsin 1.43\]

OpenStudy (anonymous):

Why did i get negative?

zepdrix (zepdrix):

Hmm I dunno, are you sure you're in `degree mode` on your calculator? It should be positive.

OpenStudy (anonymous):

okay i redid it on my other calculator and got 1.42

OpenStudy (anonymous):

im not sure why its not working on my graphing calculator

zepdrix (zepdrix):

Yah you're in radians mode on your graphing calculator :) Just switch over to degrees.

OpenStudy (anonymous):

it was in RADIANS!!! Thats why!! i hope it didnt screw up any of the other problems that ive been doing :/

zepdrix (zepdrix):

UH OH! :O

OpenStudy (anonymous):

ok so where do we go from here?

zepdrix (zepdrix):

When we try to solve for A, using the Inverse Sine Function, we get an error on our calculator. Which either meanssssss, ~It doesn't form a triangle, or, ~Maybe A is an obtuse angle. Grr I can't remember how to do these darn Law of Sines problems :p it's been so long.

OpenStudy (anonymous):

i think it doesn't form one

OpenStudy (anonymous):

How would i do one of these problems when they say that two triangles are made from whatever they give you

zepdrix (zepdrix):

Mmmmmmmmmmmmmmmmm I can't remember :C

OpenStudy (anonymous):

What about turning complex numbers into trig form?

OpenStudy (anonymous):

cause i can do it with no problem when i turn the trig form back into complex but i have a problem doing it the other way

OpenStudy (anonymous):

Well thank you for your help anyways :) Have a good night